OpenStudy (anonymous):

Write the slope-intercept equation for the line with slope 1/2 and y-intercept (0, -5)

OpenStudy (anonymous):

formula is [Y-Y1]=m[X-X1] IT SHOULD BE WRITTEN AS [Y-(-5)]=1/2[X-(0)]